A 25.00 mL solution of 0.08680 M NaI is titrated with 0.05120 M AgNO3 . Calculate pAg+ following the addition of the given volumes of AgNO3 . The ?sp of AgI is 8.3×10^−17

36.00 mLpAg+ =   ?

?e pAg+ = ?

47.40 mL pAg+ = ?
Answer #1

Consider a titration reaction as AgNO3(aq) +NaI (aq) AgI (s) + NaNO3(aq)

Calculation of equivalence point

We have, M Ag+ x V Ag+ M I-x V I-

V Ag+ M Br-x V I- / M Ag+

V Ag+    0.08680 M  x 25.00 ml / 0.05120 M

V Ag+ 42.38 ml  

Consider dissolution of AgI as AgI (s) Ag+(aq) + I -(aq)  

Ksp =[ Ag+ ][I-]=8.3 x 10-17

[ Ag+ ]=  Ksp / [I-]

1) pAg + when 36.00 ml Silver nitrate added

Before equivalence point , all added Ag+are consumed by I - and excess I - remains in the solution . We can calculate concentration of non reacted I - as

Moles of I - = original moles of I - - moles of Ag+ added

= (0.025 L ) ( 0.08680 mol/L) - (0.036 L ) ( 0.05120 mol/L) ( No. of moles = Volume in L x Molarity )

= 0.00217 mol  - 0.001843 mol

=0.000327 mol

Volume after addition of 36.00 ml AgNO3 = (0.025+0.036)= 0.061 L

  [I-] = No. of moles/ Volume of solution in L = 0.000327 mol /0.061 L = 0.00536 M

[ Ag+ ]=  Ksp / [Br-]

[ Ag+ ] = 8.3 x 10-17 /0.00536

[ Ag+ ] = 1.55 x 10-14 M

p Ag+ = -log  [ Ag+ ] = - log 1.55 x 10-14 = 13.8

2) pAg + at equivalence point

At equivalence point all Ag + is consumed by I -. But there is equilibrium between solid AgI and dissolved ions. Hence , at equilibrium the concentration of Ag + ions will be due to dissociation of solid AgI.

Consider dissolution of AgI as AgI (s)   Ag +(aq) + I -(aq)

For above reaction, Ksp =[Ag +][I - ]=8.3 10-17

If S is the solubility of AgI then [ Ag +] = [I - ]= S

Therefore, S S = 8.3 10-17

S 2 = 8.3 10-17

S = 9.11 10-09 M =[ Ag +] = [I - ]

We have, p Ag + = -log [ Ag + ]

p Ag + = -log( 9.11 10-09 )

p Ag + = 8.04

3) pAg + when 47.40 ml Silver nitrate added

After equivalence point, there is excess of Ag + ions in the solution.

Moles of Ag += added moles of Ag + - moles of I - added

Moles of Ag += (0.0474 L ) ( 0.05120 mol/L) - (0.025 L ) ( 0.08680 mol/L)

Moles of Ag += 0.00243 mol - 0.00217 mol = 0.000257 mol

Volume of solution after addition 47.40 ml of 0.05120 M Ag(NO3 )  = (0.025 + 0.0474 ) = 0.0724 L

[Ag+ ]= No. of moles / Volume of solution in L = 0.000257 mol / 0.0724 L = 0.003549 M

p Ag + = -log [Ag + ]

p Ag + = -log(0.003549)

p Ag + =2.45